# {textblock}

### v4

@@ Abstract
LLM-based intelligent agents face significant deployment challenges, particularly related to resource management. Allowing unrestricted access to LLM or tool resources can lead to inefficient or even potentially harmful resource allocation and utilization for agents. Furthermore, the absence of proper scheduling and resource management mechanisms in current agent designs hinders concurrent processing and limits overall system efficiency. To address these challenges, this paper proposes the architecture of AIOS (LLM-based AI Agent Operating System) under the context of managing LLM-based agents. It introduces a novel architecture for serving LLM-based agents by isolating resources and LLM-specific services from agent applications into an AIOS kernel. This AIOS kernel provides fundamental services (e.g., scheduling, context management, memory management, storage management, access control) for runtime agents. To enhance usability, AIOS also includes an AIOS SDK, a comprehensive suite of APIs designed for utilizing functionalities provided by the AIOS kernel. Experimental results demonstrate that using AIOS can achieve up to 2.1×2.1$\times$2.1 × faster execution for serving agents built by various agent frameworks.
[Uncaptioned image] AIOS Code Repository: https://github.com/agiresearch/AIOS
[Uncaptioned image] AIOS SDK Code Repository: https://github.com/agiresearch/Cerebrum
= Live Demo: https://app.aios.foundation
@@ 1 Introduction
In the field of autonomous agents, research efforts  (Wooldridge and Jennings, 1995, Jennings et al., 1998, Bresciani et al., 2004) are made towards agents that can perceive environments, understand instructions, make decisions, take action and learn from feedbacks. The advent of large language models (LLMs) (Achiam et al., 2023, Touvron et al., 2023a, Team et al., 2023) has brought new possibilities to the agent development Ge et al. (2023a). Current LLMs have shown great power in understanding instructions (Ouyang et al., 2022, Chung et al., 2022, Touvron et al., 2023b, Geng et al., 2022), reasoning and solving problems (Kojima et al., 2022, Nijkamp et al., 2022, Taylor et al., 2022, Hao et al., 2023, Kim et al., 2023), and interacting with human users (Ross et al., 2023) as well as external environments (Driess et al., 2023, Brohan et al., 2023). Built upon these powerful LLMs, emergent LLM-based agents (Ge et al., 2023a, Yao et al., 2023, Shinn et al., 2023, Deng et al., 2023, Packer et al., 2023, Wu et al., 2024) can present strong task fulfillment abilities in diverse environments, ranging from virtual assistants to more sophisticated reasoning and problem-solving systems.
@@
Figure 1: A motivating example of how an agent (i.e., travel agent) requires both LLM-related and Non-LLM-related (i.e., OS) services to complete a task; color in red represents services related to LLM and color in blue represents services not related to LLM.
@@ 3.4 Context Manager
LLM inference time creates bottlenecks through long-running system calls that monopolize resources. Our context interrupt mechanism addresses this via task interruption and resumption through snapshot and restoration operations. The context manager designs two approaches: text-based (for closed-source LLMs without logits access, saving decoded outputs) and logits-based (preserving intermediate search tree structure for finer-grained state restoration). The logits-based method is illustrated in Figure 4. Using beam search (common in LLMs (Touvron et al., 2023b, Jiang et al., 2023, Biderman et al., 2023)), with beam width 1 for simplicity, we demonstrate the process: When processing Determine whether there will be rain in the destination of flight UA057, the LLM evaluates candidate tokens at each step. If suspended by the scheduler mid-generation, the context manager snapshots the intermediate results. Upon resumption, it reloads this snapshot to continue from the suspension point, reaching the final answer: Search weather in Paris without restarting computation.
@@ 3.5 Memory Manager
Unlike traditional OS memory managers handling physical RAM, AIOS’s memory manager addresses agent interaction histories during runtime (Lerman and Galstyan, 2003, Zhang et al., 2024), including conversation logs and tool-calling results. It manages memory structure, allocation, read/write operations, deletion, and updates. Agent memory resides in RAM by default, but when allocated space approaches capacity, the manager implements memory swapping between RAM and disk. When an agent’s memory usage exceeds its block limit (e.g., 80% of allocation), the memory manager initiates a K-Least Recently Used (LRU-K) eviction policy, transferring items from RAM to disk via the storage manager (detailed in Section 3.6). LRU-K prioritizes retaining items in RAM that have been accessed at least K times recently, while moving less frequently accessed items to disk. This balances memory efficiency by offloading infrequently accessed data while ensuring retrievability when needed. Detailed implementations of memory manager are in A.5.
@@ 3.9 AIOS SDK
We design the AIOS SDK (Rama et al., 2025) to streamline the development and integration of agents on the AIOS architecture. This SDK not only empowers developers to build agents that interact with the core functions in the AIOS kernel but also abstracts complex system calls, allowing developers to focus on the agent’s internal workflows.
@@ 4.1 Setup
Table 1: Evaluation of agent performance on benchmarks w/o and w/ AIOS, respectively. Success rate (SR%) is used as the metric for all the benchmarks. "-" represents methods that failed GAIA benchmark tasks due to lack of API support.
@@ 4.3 Efficiency Analysis (RQ2)
In our efficiency experiments, we evaluate system performance using two key metrics: throughput and latency. Throughput is measured by counting the number of AIOS system calls executed per second, indicating the system’s capacity to handle multiple requests in parallel. Latency, on the other hand, is measured as the average waiting time experienced by agents, from the moment a query is submitted to the completion of the response, reflecting the system’s responsiveness. To ensure a controlled and consistent testing environment, we conduct these evaluations using the two open-source models, Llama-3.1-8b and Mistral-7b, both hosted locally. Hosting these models locally reduces potential variability in LLM API response times due to network-related latency issues. As shown in 6(a) and 7(a), the results demonstrate that AIOS achieves significantly higher throughput across different agent frameworks, to a 2.1×2.1$\times$2.1 × increase in throughput when using Reflexion-based agents on Llama-3.1-8b. This improvement is attributed to the scheduling employed in the AIOS kernel, which prevents unnecessary trial-and-error attempts by avoiding prompts that cannot be loaded onto the GPU for execution. In terms of latency, as illustrated in 6(b) and 7(b), the average waiting time for agents is also substantially reduced. This reduction highlights the efficiency of AIOS in serving LLM-based agents.
@@ 5 Related Work
The evolution of operating systems (OS) has progressed from rudimentary to sophisticated interactive systems. This evolution transitioned from basic batch processing (IBM, 2010) to advanced process management including time-sharing (Ritchie and Thompson, 1974) and multi-tasking (Hoare, 1974, Engler et al., 1995), enabling complex task handling. Development advanced toward modular architecture with specialized components for process scheduling (Liu and Layland, 1973, Dijkstra, 2002), memory management (Denning, 1968, Daley and Dennis, 1968), and filesystem operations (Rosenblum and Ousterhout, 1992, McKusick et al., 1984), improving system efficiency. The introduction of graphical user interfaces (GUIs) in Macintosh, Windows, and GNOME enhanced user interaction. Currently, AI models, particularly LLMs, are migrating from application to system layers, providing standardized services across applications.
Large Language Model Agents are used to solve complex planning and reasoning tasks (Xie et al., 2024, Ge et al., 2023a). Single agents engage with either digital environment or physical environment, which may invoke APIs (Ge et al., 2023a, Schick et al., 2023, Yao and Narasimhan, 2023, Parisi et al., 2022, Tang et al., 2023, Xie et al., 2024), browse websites (Nakano et al., 2022, Deng et al., 2023, Wu et al., 2024), or execute codes (Zhang et al., 2023, Yang et al., ), while agents in the physical environment may manipulate objects (Brohan et al., 2023, Fan et al., 2022, Wang et al., 2023a), carry out lab experiments (Boiko et al., 2023, Bran et al., 2023), or make actionable decisions (Huang et al., 2022, Xiang et al., 2023). LLM-based multi-agent systems (MAS) leverage the interaction among multiple agents for problem solving. The relationship among the multiple agents could be cooperative (Wang et al., 2023c, Mandi et al., 2023), competitive (Chan et al., 2023, Du et al., 2023), or a mixture of cooperation and competition (Ge et al., 2023b). In cooperative multi-agent systems, each agent takes and assesses the information provided by other agents, thereby working together to solve complex tasks, such as role playing (Li et al., 2023, Chen et al., 2023, Zhu et al., 2023), social simulation (Park et al., 2023) and software development (Hong et al., 2023, Qian et al., 2023, Wu et al., 2023, Josifoski et al., 2023).
@@ A.2 LLM Core
Backend Supported Features in AIOS Structured Output Function Calling OpenAI (cloud) $\checkmark$ $\checkmark$ Anthropic (cloud) $\checkmark$ $\checkmark$ Google (cloud) $\checkmark$ $\checkmark$ Groq (cloud) $\checkmark$ $\checkmark$ Bedrock (cloud) $\checkmark$ $\checkmark$ Huggingface (local) $\checkmark$ $\checkmark$ vllm (local) $\checkmark$ $\checkmark$ Ollama (local) $\checkmark$ $\checkmark$
@@ Appendix D Additional Experimental Results
Table 6: Impact of using different scheduling strategies, where NONE represents without using AIOS, FIFO and RR represent using AIOS with the two different scheduling strategies. All metrics are reported in seconds.

### v5

# AIOS: LLM Agent Operating System

@@ Abstract
LLM-based intelligent agents face significant deployment challenges, particularly related to resource management. Allowing unrestricted access to LLM or tool resources can lead to inefficient or even potentially harmful resource allocation and utilization for agents. Furthermore, the absence of proper scheduling and resource management mechanisms in current agent designs hinders concurrent processing and limits overall system efficiency. To address these challenges, this paper proposes the architecture of AIOS (LLM-based AI Agent Operating System) under the context of managing LLM-based agents. It introduces a novel architecture for serving LLM-based agents by isolating resources and LLM-specific services from agent applications into an AIOS kernel. This AIOS kernel provides fundamental services (e.g., scheduling, context management, memory management, storage management, access control) for runtime agents. To enhance usability, AIOS also includes an AIOS SDK, a comprehensive suite of APIs designed for utilizing functionalities provided by the AIOS kernel. Experimental results demonstrate that using AIOS can achieve up to $2.1\times$ faster execution for serving agents built by various agent frameworks. The source code is available at https://github.com/agiresearch/AIOS.
@@ 1 Introduction
In the field of autonomous agents, research efforts  (Wooldridge & Jennings, 1995; Jennings et al., 1998; Bresciani et al., 2004) are made towards agents that can perceive environments, understand instructions, make decisions, take action and learn from feedbacks. The advent of large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023a; Team et al., 2023) has brought new possibilities to the agent development Ge et al. (2023a). Current LLMs have shown great power in understanding instructions (Ouyang et al., 2022; Chung et al., 2022; Touvron et al., 2023b; Geng et al., 2022), reasoning and solving problems (Kojima et al., 2022; Nijkamp et al., 2022; Taylor et al., 2022; Hao et al., 2023; Kim et al., 2023), and interacting with human users (Ross et al., 2023) as well as external environments (Driess et al., 2023; Brohan et al., 2023). Built upon these powerful LLMs, emergent LLM-based agents (Ge et al., 2023a; Yao et al., 2023; Shinn et al., 2023; Deng et al., 2023; Packer et al., 2023; Wu et al., 2024) can present strong task fulfillment abilities in diverse environments, ranging from virtual assistants to more sophisticated reasoning and problem-solving systems.
@@
Figure 1: A motivating example of how an agent (i.e., travel agent) requires both LLM-related and Non-LLM-related (i.e., OS) services to complete a task, where color in red represents services related to LLM and color in blue represents services not related to LLM.
@@ 3.4 Context Manager
LLM inference time creates bottlenecks through long-running system calls that monopolize resources. Our context interrupt mechanism addresses this via task interruption and resumption through snapshot and restoration operations. The context manager designs two approaches: text-based (for closed-source LLMs without logits access, saving decoded outputs) and logits-based (preserving intermediate search tree structure for finer-grained state restoration). The logits-based method is illustrated in Figure 4. Using beam search (common in LLMs (Touvron et al., 2023b; Jiang et al., 2023; Biderman et al., 2023)), with beam width 1 for simplicity, we demonstrate the process: When processing Determine whether there will be rain in the destination of flight UA057, the LLM evaluates candidate tokens at each step. If suspended by the scheduler mid-generation, the context manager snapshots the intermediate results. Upon resumption, it reloads this snapshot to continue from the suspension point, reaching the final answer: Search weather in Paris without restarting computation.
@@ 3.5 Memory Manager
Unlike traditional OS memory managers handling physical RAM, AIOS’s memory manager addresses agent interaction histories during runtime (Lerman & Galstyan, 2003; Zhang et al., 2024), including conversation logs and tool-calling results. It manages memory structure, allocation, read/write operations, deletion, and updates. Agent memory resides in RAM by default, but when allocated space approaches capacity, the manager implements memory swapping between RAM and disk. When an agent’s memory usage exceeds its block limit (e.g., 80% of allocation), the memory manager initiates a K-Least Recently Used (LRU-K) eviction policy, transferring items from RAM to disk via the storage manager (detailed in Section 3.6). LRU-K prioritizes retaining items in RAM that have been accessed at least K times recently, while moving less frequently accessed items to disk. This balances memory efficiency by offloading infrequently accessed data while ensuring retrievability when needed. Detailed implementations of memory manager are in A.5.
@@ 3.9 AIOS SDK
We design the AIOS SDK to streamline the development and integration of agents on the AIOS architecture. This SDK not only empowers developers to build agents that interact with the core functions in the AIOS kernel but also abstracts complex system calls, allowing developers to focus on the agent’s internal workflows.
@@ 4.1 Setup
Table 1: Evaluation of agent performance on benchmarks w/o and w/ AIOS, respectively. Success rate (SR%) is used as the metric for all the benchmarks. ”-” represents methods that failed GAIA benchmark tasks due to lack of API support.
@@ 4.3 Efficiency Analysis (RQ2)
In our efficiency experiments, we evaluate system performance using two key metrics: throughput and latency. Throughput is measured by counting the number of AIOS system calls executed per second, indicating the system’s capacity to handle multiple requests in parallel. Latency, on the other hand, is measured as the average waiting time experienced by agents, from the moment a query is submitted to the completion of the response, reflecting the system’s responsiveness. To ensure a controlled and consistent testing environment, we conduct these evaluations using the two open-source models, Llama-3.1-8b and Mistral-7b, both hosted locally. Hosting these models locally reduces potential variability in LLM API response times due to network-related latency issues. As shown in 6(a) and 7(a), the results demonstrate that AIOS achieves significantly higher throughput across different agent frameworks, to a $2.1\times$ increase in throughput when using Reflexion-based agents on Llama-3.1-8b. This improvement is attributed to the scheduling employed in the AIOS kernel, which prevents unnecessary trial-and-error attempts by avoiding prompts that cannot be loaded onto the GPU for execution. In terms of latency, as illustrated in 6(b) and 7(b), the average waiting time for agents is also substantially reduced. This reduction highlights the efficiency of AIOS in serving LLM-based agents.
@@ 5 Related Work
The evolution of operating systems (OS) has progressed from rudimentary to sophisticated interactive systems. This evolution transitioned from basic batch processing (IBM, 2010) to advanced process management including time-sharing (Ritchie & Thompson, 1974) and multi-tasking (Hoare, 1974; Engler et al., 1995), enabling complex task handling. Development advanced toward modular architecture with specialized components for process scheduling (Liu & Layland, 1973; Dijkstra, 2002), memory management (Denning, 1968; Daley & Dennis, 1968), and filesystem operations (Rosenblum & Ousterhout, 1992; McKusick et al., 1984), improving system efficiency. The introduction of graphical user interfaces (GUIs) in Macintosh, Windows, and GNOME enhanced user interaction. Currently, AI models, particularly LLMs, are migrating from application to system layers, providing standardized services across applications.
Large Language Model Agents are used to solve complex planning and reasoning tasks (Xie et al., 2024; Ge et al., 2023a). Single agents engage with either digital environment or physical environment, which may invoke APIs (Ge et al., 2023a; Schick et al., 2023; Yao & Narasimhan, 2023; Parisi et al., 2022; Tang et al., 2023; Xie et al., 2024), browse websites (Nakano et al., 2022; Deng et al., 2023; Wu et al., 2024), or execute codes (Zhang et al., 2023; Yang et al., ), while agents in the physical environment may manipulate objects (Brohan et al., 2023; Fan et al., 2022; Wang et al., 2023a), carry out lab experiments (Boiko et al., 2023; Bran et al., 2023), or make actionable decisions (Huang et al., 2022; Xiang et al., 2023). LLM-based multi-agent systems (MAS) leverage the interaction among multiple agents for problem solving. The relationship among the multiple agents could be cooperative (Wang et al., 2023c; Mandi et al., 2023), competitive (Chan et al., 2023; Du et al., 2023), or a mixture of cooperation and competition (Ge et al., 2023b). In cooperative multi-agent systems, each agent takes and assesses the information provided by other agents, thereby working together to solve complex tasks, such as role playing (Li et al., 2023; Chen et al., 2023; Zhu et al., 2023), social simulation (Park et al., 2023) and software development (Hong et al., 2023; Qian et al., 2023; Wu et al., 2023; Josifoski et al., 2023).
@@ 6 Conclusion and Future Work
7 Acknowledgement
We thank Balaji Rama, Hang Gao, Shuhang Lin, Jian Zhang and Zhenting Wang for their valuable discussions and suggestions during the project.
@@ A.2 LLM Core
Backend Supported Features in AIOS Structured Output Function Calling OpenAI (cloud) Anthropic (cloud) Google (cloud) Groq (cloud) Bedrock (cloud) Huggingface (local) vllm (local) Ollama (local)
@@ Appendix D Additional Experimental Results
Table 6: Impact of using different scheduling strategies, where NONE represents without using AIOS, FIFO and $R$ represent using AIOS with the two different scheduling strategies. All metrics are reported in seconds.